# How Does Knowledge Selection Help Retrieval Augmented Generation?

### v2

@@ Abstract
Retrieval-augmented generation (RAG) is a powerful method for enhancing natural language generation by integrating external knowledge into a model’s output. While prior work has demonstrated the importance of improving knowledge retrieval for boosting generation quality, the role of knowledge selection remains less clear. In this paper, we perform a comprehensive analysis of how knowledge selection influences downstream generation performance in RAG systems. By simulating different retrieval and selection conditions through a controlled mixture of gold and distractor knowledge, we assess the impact of these factors on generation outcomes. Our findings indicate that the downstream generator model’s capability, as well as the complexity of the task and dataset, significantly influence the impact of knowledge selection on the overall RAG system performance. In typical scenarios, improving the knowledge recall score is key to enhancing generation outcomes, with the knowledge selector providing a limited additional benefit when a strong generator model is used on clear, well-defined tasks. For weaker generator models or more ambiguous tasks and datasets, the knowledge F1 score becomes a critical factor, and the knowledge selector plays a more prominent role in improving overall performance.
@@ 1 Introduction
The quality of the retrieved knowledge directly influences the quality of RAG outputs. Previous studies have consistently shown that improving knowledge retrieval leads to a direct enhancement in generation performance Dinan et al. (2019); Li et al. (2022, 2024); Cuconasu et al. (2024); Wang et al. (2024); Wu et al. (2024). Similarly, effective knowledge selection has been observed to improve generation quality by filtering out retrieved information that is less relevant to the generation target Kim et al. (2020); Thulke et al. (2021); Li et al. (2022); Sun et al. (2023); Zhang et al. (2023); Zheng et al. (2024); Wang et al. (2024). However, we observe that knowledge selection is rarely used for tasks other than dialogue generation, and there are notably fewer LLM-based RAG works that use knowledge selection, compared to fine-tuned RAG models (see Section 2). We hypothesize that knowledge selectors may not always improve downstream generation performance, and that there may be a selection bias where only positive results involving knowledge selector modules are published, while experiments where knowledge selectors are not helpful simply do not report those results.
In this work, we perform a systematic analysis of how knowledge selectors with various performances affect downstream RAG performance. By blending gold knowledge with distractor knowledge in varying ratios, we simulate different selection outcomes and examine their impact on the overall performance of RAG. We find that the generator model’s capability, as well as the complexity of the task and dataset, significantly influence RAG system performance. In typical scenarios, improving knowledge recall via the knowledge retriever is key to enhancing generation outcomes, with the knowledge selector providing limited additional benefit when a strong generator model is used on clear, well-defined tasks. For weaker generator models or more ambiguous tasks and datasets, knowledge F1 becomes a critical factor, and the knowledge selector plays a more prominent role in improving overall performance.
Figure 1: Our simulation experiment pipeline for each example.
@@ 2 Related Work
Dialogue generation Moghe et al. (2018); Dinan et al. (2019); Li et al. (2024) is one of the major applications of RAG, where the target response is conditioned on retrieved knowledge, and a knowledge selection step is often added to further refine the retrieved knowledge Thulke et al. (2021); Sun et al. (2023); Zhang et al. (2023). For example, Kim et al. (2020) train a knowledge selector by leveraging response information, Li et al. (2022) select knowledge from document semantic graphs, and Zhang et al. (2023) propose multi-task learning for knowledge selection and response generation. However, despite these works demonstrating the advantages of their overall approaches through ablation studies, it is unclear whether their performance improvement via knowledge selection specifically is transferable to other datasets, domains, or tasks. Moreover, despite the popularity of LLMs for RAG, there are dramatically fewer LLM-based RAG works that include the knowledge selection step, and it is likewise omitted in other RAG tasks Gao et al. (2023); Fan et al. (2024).
To the best of our knowledge, there is no prior work explaining this gap in knowledge selection. The closest work to ours are Cuconasu et al. (2024) and Wu et al. (2024), which study the impact of knowledge retrieval on RAG. We hypothesize that there is a selection bias effect where knowledge selection modules may have been experimented with, but ultimately not reported, in tasks or settings where it is not helpful. In this paper, we study the general effect of knowledge retrieval and selection on LLM-based RAG via simulations, rather than being restricted to specific retriever or selector implementations.
@@ 3.1 Task Formulation
RAG involves three steps: (1) Knowledge retrieval, where a retriever module retrieves a set of candidate knowledge $K$ based on a query $q$. This step aims to retrieve as much knowledge that is relevant to the query as possible, attempting to balance knowledge recall and precision. (2) Optionally, knowledge selection removes retrieved knowledge that is less relevant to further improve the knowledge precision, producing $K^{\prime}\subseteq K$. (3) Finally in the generation step, the generator takes the query $q$ and the selected retrieved knowledge $K^{\prime}$ to generate the output text $r$.
The knowledge in $K$ and $K^{\prime}$ can be heterogeneous and come from multiple sources, such as external documents, knowledge graphs, or conversation histories. This RAG framework can be applied to various tasks, such as dialogue generation, question answering, fact verification, and code generation Gao et al. (2023); Fan et al. (2024). In our experiments on dialogue generation and question answering, we simulate steps 1 and 2 by creating controlled knowledge sets $K^{\prime}$ and observing the resulting performance of step 3.
@@ 3.2 Knowledge Simulation
We aim to perform a systematic analysis of the effect of knowledge retrieval and selection outcomes on the downstream RAG performance by simulation. As Figure 1 shows, for each query $q$, given a fixed pool of available knowledge with gold relevance annotations, we sample gold knowledge and distractor knowledge at varying rates $p_{gold}$ and $p_{noise}$ to precisely simulate a wide range of quality for the retrieved and selected knowledge $K^{\prime}$, from noise only to gold knowledge only. For example, if $p_{gold}=0.5$, then each piece of gold knowledge has a 50% chance to be sampled. In this way, we simulate a wide distribution of knowledge retriever and selector performance, as measured by knowledge precision and recall computed based on the annotated gold knowledge. We then test the performance of the generator model given the different quality of simulated retrieved and selected knowledge $K^{\prime}$.
@@ 4.1 Datasets
While there are several datasets supporting the RAG framework, relatively few provide high-quality, human-annotated gold knowledge. We use two popular and representative datasets, Wizard of Wikipedia (WoW; Dinan et al., 2019) and HotpotQA Yang et al. (2018) for our experiments.
@@ 4.2 Generators
To align with the latest trends in RAG research, as well as to make the analysis simple and generally replicable, we adopt LLMs as our generators. We use three API-based LLMs, OpenAI GPT-4o-mini, LLaMA 3.1 8B, and Mistral 7B-Instruct, as our generator models; the varying performances of the LLMs allow us to investigate the impact of generator complexity on knowledge usage.
@@ 4.3 Knowledge Sampling
As Figure 1 shows, both WoW and HotpotQA are knowledge selector training datasets that provide a retrieved knowledge set $K$ for each query $q$ such that the gold knowledge is a subset of $K$. In both datasets, $K$ contains text passages consisting of an article title and a few knowledge sentences. We perform knowledge sampling at the sentence level to simulate the end result of applying both the knowledge retrieval and selection steps.
In our experiments, we refer to using the entire retrieved knowledge set $K$ (provided by the dataset) as the input knowledge to the generator (i.e. no knowledge selection) as the “full knowledge” setting, where both gold and distractor knowledge are present. We also include the “no knowledge" setting, where the generators receive no external knowledge at all, as a weak baseline, and “gold knowledge," corresponding to perfect knowledge selection, as an upper bound.
Figure 2: Scatter plot of response/answer F1, plotted against knowledge precision (x-axis) and recall (y-axis), by GPT-4o-mini (top), LLaMA 3.1 8B (middle), and Mistral 7B-Instruct (bottom). The left column shows results on WoW; the right shows HotpotQA. The dots highlighted in orange indicate settings outperforming the “full knowledge” setting, while those highlighted in cyan indicate settings underperforming the “no knowledge" setting. Each dot corresponds to a full experiment on the entire sampled dataset.
5 Simulation Results
Overall we observe several consistent trends, except for Mistral-7B-Instruct, which is a relatively weak-performing LLM. Note that each point plotted in Figures 2-4 corresponds to a full experiment on the entire sampled dataset. The standard error mean values in Tables 1 & 2 indicate that the answer F1 scores are robust across all experiments and the overall observed trends are significant.
@@ 5 Simulation Results
As Tables 1 & 2 show, generators without any retrieved knowledge (“no knowledge" setting) perform poorly on WoW and HotpotQA, even though they may have been pre-trained on the Wikipedia articles that WoW and HotpotQA are derived from. This finding indicates the LLMs are not over-fitted on the WoW and HotpotQA datasets, and applying RAG is beneficial.
Interestingly, our HotpotQA results in Figure 2 show that distractor knowledge significantly harms performance; the cyan dots show that generators receiving mostly distractor knowledge underperform the “no knowledge” setting, which uses only the LLM’s internal knowledge. This trend is not observed with WoW, supporting our observation that HotpotQA’s distractor knowledge is truly irrelevant, while the “distractor" knowledge in WoW may only be less relevant than the single gold sentence, but not completely irrelevant. This difference between the two datasets is further explored in Appendix A.4.
Figure 3: Scatter plot of HotpotQA answer F1 versus knowledge precision (top), knowledge recall (middle), and knowledge F1 (bottom). The left column shows GPT-4o-mini as the generator; the right column shows Mistral-7B-Instruct. Plots for LLaMa 3.1 8B and the WoW dataset are in Appendix A. Each point corresponds to a full experiment on the entire sampled dataset.
@@
Figure 4 further shows that increases in knowledge recall correspond to significant increases in answer F1 scores. Taking GPT-4o-mini on HotpotQA as an example, we see that moving from the left end of a color contour to the right (i.e. keeping knowledge recall fixed while improving precision) only slightly improves answer F1. In contrast, moving from one contour to another (i.e. varying knowledge recall) significantly impacts answer F1 for non-zero precision scores. Thus, to improve knowledge quality for RAG, improving the retriever’s recall score is the top priority; improving precision via a knowledge selector has a limited contribution to generation performance, especially for strong generators.
For weaker generators, like Mistral 7B-Instruct, the relationship between generation and knowledge F1 is stronger, while the correlation with recall is weaker (Figure 3, right column). Similarly, we see much less separation between color contours (recall) and a much steeper increase in answer F1 with precision (Figure 4, bottom row).
@@
All of our results show that the overall RAG performance depends on the generator model: the cross-model comparisons in Figures 2 & 4 show that for stronger generator models (as measured by rank in LLM leaderboards***https://huggingface.co/spaces/lmsys/chatbot-arena-leaderboard), performance across all knowledge settings are higher (Tables 1 & 2). While this result is unsurprising, we also find that the gap between the “full knowledge” and “gold knowledge” settings becomes narrower, suggesting that stronger generators are more robust to noisy input knowledge and rely less on the knowledge selector. In contrast, when the generator is weaker, any reasonable knowledge selector becomes beneficial, likely because a weak generator cannot handle noisy input and requires a selector to filter out distractor knowledge — knowledge F1 best correlates with answer F1 for Mistral-7B-Instruct in Figure 3.
@@
In addition, Figure 4 shows that attempting to improve the knowledge selector on top of a weak retriever (i.e. increasing precision when recall is low) hurts generation performance on WoW, as well as for GPT-4o-mini on Hotpot QA. In this scenario, more total knowledge, regardless of noise, improves response generation. These counter-intuitive observations are likely due to the nature of the task and the annotation quality. The solution space in HotpotQA is small given a specific question, and HotpotQA has a cleaner separation between gold and distractor knowledge, whereas in WoW, there are more plausible responses given the same conversation history and knowledge, as well as “distractor" knowledge that may actually be relevant to the gold response (see Appendix A.4 for more analysis).
@@
Interestingly, we observe that the boundary of where the knowledge selector improves generation performance (the border between the orange and white areas in Figure 2) is convex for all three LLMs on WoW. This phenomenon can also be seen in Figure 4, where some color contours intersect with the “full knowledge answer F1” baseline (dashed red line) multiple times. In other words, we observe that generation performance is non-monotonic as knowledge precision increases with a fixed knowledge recall. The fact that this phenomenon is only observed on WoW may be due to the relatively noisy gold knowledge annotations in WoW; we can produce similar behavior by artificially injecting noise into HotpotQA’s gold knowledge annotations (Appendix A.5).
@@
One important motivation for using a knowledge selector is to reduce the total input length to the generator model. However, while computational costs can be reduced by only using the top-$k$ knowledge, we find that the overall relationship between the knowledge precision-recall and generation F1 observed in our simulations remains unchanged, regardless of the value of $k$ (Appendix A.6).

### v3

@@ Abstract
Retrieval-augmented generation (RAG) is a powerful method for enhancing natural language generation by integrating external knowledge into a model’s output. While prior work has demonstrated the importance of improving knowledge retrieval for boosting generation quality, the role of knowledge selection remains less clear. This paper empirically analyzes how knowledge selection influences downstream generation performance in RAG systems. By simulating different retrieval and selection conditions through a controlled mixture of gold and distractor knowledge, we assess the impact of these factors on generation outcomes. Our findings indicate that the downstream generator model’s capability, as well as the complexity of the task and dataset, significantly influence the impact of knowledge selection on the overall RAG system performance. In typical scenarios, improving the knowledge recall score is key to enhancing generation outcomes, with the knowledge selector providing limited benefit when a strong generator model is used on clear, well-defined tasks. For weaker generator models or more ambiguous tasks and datasets, the knowledge F1 score becomes a critical factor, and the knowledge selector plays a more prominent role in improving overall performance.
@@ 1 Introduction
The quality of the retrieved knowledge directly influences the quality of RAG outputs. Previous studies have consistently shown that improving knowledge retrieval leads to a direct enhancement in generation performance Dinan et al. (2019); Li et al. (2022, 2024); Cuconasu et al. (2024); Wang et al. (2024); Wu et al. (2024). Similarly, effective knowledge selection has been observed to improve generation quality by filtering out retrieved information that is less relevant to the generation target Kim et al. (2020); Thulke et al. (2021); Li et al. (2022); Sun et al. (2023); Zhang et al. (2023); Zheng et al. (2024); Wang et al. (2024). However, we observe that knowledge selection is rarely used for tasks other than dialogue generation, and there are notably fewer LLM-based RAG works that use knowledge selection, compared to fine-tuned RAG models (see Section 2).
We hypothesize that knowledge selectors may not always improve downstream generation performance, and that there may be a selection bias where only positive results involving knowledge selector modules are published, while experiments where knowledge selectors are not helpful simply do not report those results. Further, prior works focus on proposing specific knowledge selection approaches, offering only narrow, case-specific insights into the relationship between knowledge and generation. As a result, readers are often left with anecdotal observations, such as “using model A in scenario X improves performance”, without a global picture of how knowledge selection impacts generation and when it is most effective.
Therefore, in this work, we perform a systematic empirical analysis of how knowledge selectors with various performances affect downstream RAG performance. By blending gold knowledge with distractor knowledge in varying ratios, we simulate different selection outcomes and examine their impact on the overall performance of RAG. We find that the generator model’s capability, as well as the complexity of the task and dataset, significantly influence RAG system performance. In typical scenarios, improving knowledge recall via the knowledge retriever is key to enhancing generation outcomes, with the knowledge selector providing limited additional benefit when a strong generator model is used on clear, well-defined tasks. For weaker generator models or more ambiguous tasks and datasets, knowledge F1 becomes a critical factor, and the knowledge selector plays a more prominent role in improving overall performance.
Figure 1: Our simulation experiment pipeline.
@@ 2 Related Work
Dialogue generation Moghe et al. (2018); Dinan et al. (2019); Li et al. (2024) is one of the major applications of RAG, where the target response is conditioned on retrieved knowledge, and a knowledge selection step is often added to further refine the retrieved knowledge Thulke et al. (2021); Sun et al. (2023); Zhang et al. (2023). For example, Kim et al. (2020) train a knowledge selector by leveraging response information, Li et al. (2022) select knowledge from document semantic graphs, and Zhang et al. (2023) propose multi-task learning for knowledge selection and response generation. However, while these works demonstrate the advantages of their approaches through ablation studies, it is unclear whether their performance improvement via knowledge selection specifically is transferable to other datasets, domains, or tasks. Moreover, despite the popularity of LLMs for RAG, there are dramatically fewer LLM-based works that include the knowledge selection step, and it is likewise omitted in other RAG tasks Gao et al. (2023); Fan et al. (2024).
To our knowledge, there is no prior work explaining this gap in knowledge selection. The closest works to ours are Cuconasu et al. (2024) and Wu et al. (2024), which study the impact of knowledge retrieval on RAG. We hypothesize that there is a selection bias effect where knowledge selection modules may have been experimented with, but ultimately not reported, in tasks or settings where it is not helpful. In this paper, we study the general effect of knowledge retrieval and selection on LLM-based RAG via hundreds of simulations, rather than being restricted to a few specific, anecdotal retrievers or selector implementations like prior works.
@@ 3.1 Task Formulation
RAG involves three steps: (1) Knowledge retrieval, where a retriever module retrieves a set of candidate knowledge $K$ based on a query $q$. This step aims to retrieve as much knowledge that is relevant to the query as possible, balancing knowledge recall and precision. (2) Optionally, knowledge selection removes retrieved knowledge that is less relevant to further improve knowledge precision, producing $K^{\prime}\subseteq K$. (3) Finally, the generator takes the query $q$ and the selected, retrieved knowledge $K^{\prime}$ to generate the output text $r$.
The knowledge in $K$ and $K^{\prime}$ can be heterogeneous and come from multiple sources, such as external documents, knowledge graphs, or conversation histories. This RAG framework can be applied to various tasks, such as dialogue generation, question answering, fact verification, and code generation Gao et al. (2023); Fan et al. (2024). In our experiments on dialogue generation and question answering, we simulate steps 1 and 2 by creating controlled knowledge sets $K^{\prime}$ and observing the resulting generation performance of step 3.
@@ 3.2 Knowledge Simulation
We aim to perform a systematic analysis of the effect of knowledge retrieval and selection outcomes on the downstream RAG performance by simulation. As Figure 1 shows, for each query $q$, given a fixed pool of available knowledge with gold relevance annotations, we sample gold knowledge and distractor knowledge at varying rates $p_{gold}$ and $p_{noise}$ to precisely simulate a wide range of quality for the retrieved and selected knowledge $K^{\prime}$, from noise only to gold knowledge only. For example, if $p_{gold}=0.5$, then each piece of gold knowledge has a 50% chance to be sampled. Each sampling produces a full experiment over the entire test set, reported as one data point in Figures 2-4.
In this way, we simulate a wide distribution of knowledge retriever and selector performance, as measured by knowledge precision and recall based on the gold annotations. We then test the performance of the generator model given the different quality of simulated retrieved and selected knowledge $K^{\prime}$. Compared to prior works discussed in Section 2, which only compare a few ablation configurations, we conduct hundreds of configurations in each meta-experiment.
@@ 4.1 Datasets
While there are several datasets supporting the RAG framework, relatively few provide high-quality, human-annotated gold knowledge. Furthermore, the target outputs should be relatively short, unambiguous, and easily evaluated with automatic metrics such as F1 scores. These conditions narrow down the choices to two popular and representative datasets, Wizard of Wikipedia (WoW; Dinan et al., 2019) and HotpotQA Yang et al. (2018).
@@ 4.2 Generators
To align with the latest trends in RAG research, as well as to make the analysis simple and generally replicable, we adopt LLMs as our generators. Due to the large computational costs for the meta-experiments, each of which consists of hundreds of full experiments, we use three API-based lightweight LLMs, OpenAI GPT-4o-mini, LLaMA 3.1 8B, and Mistral 7B-Instruct, as our generator models; the varying performances of the LLMs allow us to investigate the impact of generator complexity on knowledge usage.
@@ 4.3 Knowledge Sampling
As Figure 1 shows, both WoW and HotpotQA are knowledge selector training datasets that provide a retrieved knowledge set $K$ for each query $q$ such that the gold knowledge is a subset of $K$. In both datasets, $K$ contains text passages consisting of an article title and a few knowledge sentences. We perform knowledge sampling at the sentence level to simulate the end result of applying both the knowledge retrieval and selection steps. More details are in Appendix A.1.1.
In our experiments, we refer to using the entire retrieved knowledge set $K$ provided by the dataset as the input knowledge to the generator (i.e. no knowledge selection) as the “full knowledge” setting, where both gold and distractor knowledge are present. We also include a “no knowledge" setting, where the generators receive no external knowledge at all, as a weak baseline, and “gold knowledge," corresponding to perfect knowledge selection, as an upper bound.
Figure 2: Scatter plot of response/answer F1, plotted against knowledge precision (x-axis) and recall (y-axis), by GPT-4o-mini (top), LLaMA 3.1 8B (middle), and Mistral 7B-Instruct (bottom). The left column shows results on WoW; the right shows HotpotQA. The dots highlighted in orange indicate settings outperforming the “full knowledge” setting, while those highlighted in cyan indicate settings underperforming the “no knowledge" setting. Each figure is a meta-experiment, and each data point corresponds to a full experiment on the entire sampled dataset.
5 Meta-Experimental Results
Overall we observe several consistent trends across meta-experiments, except for Mistral-7B-Instruct, which is a relatively weak-performing LLM. Note that each of the plots in Figures 2-4 corresponds to a meta-experiment, and each point corresponds to a full experiment on the entire test set. The standard error mean values in Tables 1 & 2 indicate that the answer F1 scores are robust across all experiments and the overall observed trends are significant.
@@ 5 Meta-Experimental Results
As Tables 1 & 2 show, generators without any retrieved knowledge (“no knowledge") perform poorly on WoW and HotpotQA, even though they may have been pre-trained on the Wikipedia articles that WoW and HotpotQA are derived from. This finding indicates the LLMs are not over-fitted on the WoW and HotpotQA datasets, and applying RAG is beneficial.
Interestingly, our HotpotQA results in Figure 2 show that distractor knowledge significantly harms performance; the cyan dots show that generators receiving mostly distractor knowledge underperform the “no knowledge” setting, which uses only the LLM’s internal knowledge. This trend is not observed with WoW, supporting our observation that HotpotQA’s distractor knowledge is truly irrelevant, while the “distractor" knowledge in WoW may only be less relevant than the single gold sentence, but not completely irrelevant. This difference between datasets is further explored in Appendix A.4.
Figure 3: Scatter plot of HotpotQA answer F1 versus knowledge precision (top), knowledge recall (middle), and knowledge F1 (bottom). The left column shows GPT-4o-mini as the generator; the right column shows Mistral-7B-Instruct. Plots for LLaMa 3.1 8B and the WoW dataset are in Appendix A. Each figure is a meta-experiment, and each data point corresponds to a full experiment on the entire sampled dataset.
@@
Figure 4 further shows that increases in knowledge recall correspond to significant increases in answer F1 scores. Taking GPT-4o-mini on HotpotQA as an example, we see that moving from the left end of a color contour to the right (i.e. keeping knowledge recall fixed while improving precision) only slightly improves answer F1. In contrast, moving from one contour to another (i.e. varying knowledge recall) significantly impacts answer F1 for non-zero precision scores. Thus, to improve knowledge quality for RAG, improving the retriever’s recall score is the top priority; improving precision via a knowledge selector has a limited contribution, especially for strong generators.
For weaker generators, like Mistral 7B-Instruct, the relationship between generation and knowledge F1 is stronger, and correlation with recall is weaker (Figure 3 right). We see much less separation between recall color contours and a much steeper increase with precision (Figure 4 bottom).
@@
All of our results show that the overall RAG performance depends on the generator model: the cross-model comparisons in Figures 2 & 4 show that for stronger generator models (as measured by rank in LLM leaderboards [^1] https://huggingface.co/spaces/lmsys/chatbot-arena-leaderboard), performance across all knowledge settings are higher (Tables 1 & 2). While this result is unsurprising, we also find that the gap between the “full knowledge” and “gold knowledge” settings becomes narrower, suggesting that stronger generators are more robust to noisy input knowledge and rely less on the knowledge selector. In contrast, when the generator is weaker, any reasonable knowledge selector becomes beneficial, likely because a weak generator cannot handle noisy input and requires a selector to filter out distractor knowledge; knowledge F1 best correlates with answer F1 for Mistral-7B-Instruct in Figure 3.
@@
In addition, Figure 4 shows that attempting to improve the knowledge selector on top of a weak retriever (i.e. increasing precision when recall is low) hurts generation performance on WoW, as well as GPT-4o-mini on Hotpot QA. In this scenario, more total knowledge, regardless of noise, improves response generation. These counter-intuitive observations are likely due to the nature of the task and the annotation quality. The solution space in HotpotQA is small given a specific question, and HotpotQA has a cleaner separation between gold and distractor knowledge, whereas in WoW, there are more plausible responses given the same conversation history and knowledge, as well as “distractor" knowledge that may actually be relevant to the gold response (see Appendix A.4 for more analysis).
@@
Interestingly, we observe that the boundary of where the knowledge selector improves generation performance (the border between the orange and white areas in Figure 2) is convex for all three LLMs on WoW. This phenomenon can also be seen in Figure 4, where some color contours intersect with the “full knowledge” baseline (dashed red line) multiple times. In other words, we observe that generation performance is non-monotonic as knowledge precision increases with a fixed knowledge recall. The fact that this phenomenon is only observed on WoW may be due to the relatively noisy gold knowledge annotations in WoW; we can produce similar behavior by artificially injecting noise into HotpotQA’s gold knowledge annotations (Appendix A.5).
@@
One important motivation for using a knowledge selector is to reduce the total input length to the generator model. However, while computational costs can be reduced by only using the top-$k$ knowledge, we find that the overall relationship between knowledge precision-recall and generation F1 observed in our simulations remains unchanged, regardless of the value of $k$ (Appendix A.6).